non passing test has been recording the number of times his students have been previously non passing the text X for either passing and not. that X had the following probability

X:0,1,2,3

probability: 0.62, 0.17, 0.08, 0.13

a)get the probability that randomly chosen individual has never fail a test

b)get the probability for who non passing the test for the first time

c)explain what P(X<=2) means in plain language and calculate P(X<=2)
Answer #1

( a ) P (   never fail a test ) = P ( X =0 ) = 0.62

( b ) P ( non passing the test for the first time ) = P ( X = 1 ) = 0.17

( c ) P ( x <= 2 ) means that non passing the test for fewer than two times ( or ) atmost 2 times

P ( x <= 2 ) = P ( ( X =0 ) + P ( X = 1 ) + P ( X = 2 )

= 0.62 + 0.17 +  0.08

= 0.87